OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3 @satellite73

OpenStudy (anonymous):

@hartnn @Luis_Rivera @waterineyes

OpenStudy (jim766):

you would use graphing if both equations start with y = you would use substution if at least one of the equations had a variable alone do you think you should use one of these?

OpenStudy (anonymous):

I think...

OpenStudy (jim766):

I say neither of those fit... elimination is best used when the equations are in standard form...like these are. Chose elimimination...

OpenStudy (jim766):

do you know how to do it?

OpenStudy (anonymous):

I don't. I'm not good at these at all. And I don't know what you're trying to explain. @hartnn do you know what he's trying to explain because I don't understand.

hartnn (hartnn):

yes.

OpenStudy (jim766):

We have 2 variables, we need to get rid of one of them, 4x +3y = -1 3x + y = 3 )3 if I mult this eq by -3, and then add them togther the 4x + 3y = -1 y's will be eliminated -9x -3y = -9 _____________ -5x = -10 now this is an equation I can solve

OpenStudy (jim766):

x = 2, now do you know how to figure out what y is?

OpenStudy (anonymous):

Plug in x=2 to find y?

OpenStudy (jim766):

yes, it doesnt matter which one, which ever looks easiet

OpenStudy (jim766):

3(2) + y = 3 6 + y = 3 -6 -6 y = -3 x y the point were these two line intersect is ( 2, -3)

OpenStudy (jim766):

are we good?

OpenStudy (anonymous):

Kind of. How would I describe that altogether?

OpenStudy (anonymous):

Sorry I am late...

OpenStudy (anonymous):

It's okay @waterineyes

OpenStudy (anonymous):

So nice of you...

OpenStudy (jim766):

Say that you used elimination, because the equations were in standard form, and you could see it wouldnt be too hard to eliminate one.

OpenStudy (jim766):

* eliminate one variable

OpenStudy (jim766):

do you have to describe how you solve it too?

OpenStudy (anonymous):

Here I think we can use Substitution Method also efficiently because y is alone there...

OpenStudy (anonymous):

Never Mind... Sometimes, I speak uselessly..

OpenStudy (anonymous):

yes @Jim766 I have to describe how to solve it

OpenStudy (jim766):

say you multiplied the 2nd equation by -3, then added both equations together.. with the y's eliminated, you solve the equation for x. Then you substituted the value of x into the 2nd equation to find y

OpenStudy (jim766):

take a look, isnt that what we did?

OpenStudy (anonymous):

Yes! Thank you!